iamhayley:

Determine the slope and y-intercept of the line. \(y=\dfrac43x-10.7\) a. Slope =\(\dfrac43x\) y-intercept is (0, 10.7) b. Slope =\(\dfrac43\) y-intercept is (0, 10.7) c. Slope =\(\dfrac43\) y-intercept is (0, -10.7) d. Slope = -10.7 y-intercept is \(\dfrac43\)

Thera:

This in slope-intercept form. Compare [tex]y=\frac{4}{3}x-10.7[/tex] to [tex]y=mx+b[/tex]. [tex]m[/tex] is the slope. [tex]b[/tex] is the y-intercept. (The y-intercept as an ordered pair is (0,b).) Example: The slope of [tex]y=-\frac{5}{7}x+10.5[/tex] is [tex]-\frac{5}{7}[/tex]. The y-intercept of [tex]y=-\frac{5}{7}x+10.5[/tex] is 10.5. The y-intercept as an ordered pair is (0,10.5). So what do you think the slope (the m value) and the y-intercept (the b value is) of [tex]y=\frac{4}{3}x-10.7[/tex]. Let me know your thoughts.